The ages of two friends Ani and Biju differ by 3 years. Ani’s father Dharam is twice as old as
Ani and Biju is twice as old as his sister Cathy. The ages of Cathy and Dharam differ by 30
years. Find the ages of Ani and Biju.

Grade:10

3 Answers

venkatesh vadde
10 years ago

Ani - 19 years

Buju - 16 years

Cathy - 8 years

Dharam - 38 years

jitender lakhanpal
10 years ago

hi shreesha

let the ages of ani and biju be x and y

so mod of x - y = 3----1  (as we dont know who is elder thats why mod)

let age of Dharam be z and Cathy be w

z = 2x-----2   ,    y = 2 w---------3

z - w = 30 -------4

so there are 4 equations and 4 variables you can easily solve them

jitender

Sneha
5 years ago
Let Ani age = x
Bizu age=Y 
Dharma age=a
Cathy age=b
x-y=3_.......(1)
a=2x.......(2)
2b=Y......(3)
a-b=3 (given)
Put eq 2 nd 3
2x-y/2= 30
4x-y=60....(4)
 
FROM eq 1 and 4 
4(3+y) - y =60
12+4y-y= 60
Y (biju)=16 yrs
 Put y value in eq 1
 
x-16=3
X(Ali)=19 yrs
Put x and y in 2 and 3 eq 
We get
a(dharma)=38 years
B( cathy) =8 years
 
 
X= 19